# Understanding the Ranking Loss for Recommendation with Sparse User Feedback

### v1

@@ 3.2. Gradients Analysis
First, we analyze the gradients of the BCE method. Based on the chain rule, the gradients of parameters in each layer are influenced by the gradient of the logits. Therefore, we first discussed the gradients concerning the logits.
3.2.1. Gradient of BCE Loss for Negative Sample
@@
$\nabla_{z^{(-)}_{j}}\mathcal{L}_{\text{BCE}}=$ $\frac{1}{N}\cdot\frac{1}{1-\sigma(z^{(-)}_{j})}\cdot\sigma(z^{(-)}_{j})(1-\sigma(z^{(-)}_{j}))$
(9) $=$ $\frac{1}{N}\cdot\sigma(z^{(-)}_{j})=\frac{1}{N}\cdot{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}\hat{p}_{j}}.$
This equation demonstrates that the gradient of the negative sample is proportional to its pCTR value, $\hat{p}_{j}$. The expected value of $\hat{p}_{j}$ produced by an unbiased CTR estimation model with BCE loss is close to the underlying global CTR, which latter represents the proportion of click samples (i.e., positive feedback) to the total samples. This is because the BCE loss function is scale calibrated (Yan et al., 2022) and its global minima is achieved at $\sigma(z)\rightarrow E[y|x]$ (i.e., globally unbiased) (Bai et al., 2023).
When the positive feedback is sparse, $\hat{p_{j}}$ becomes a small value. According to Eq. 3.2.1, the gradient of negative samples is proportional to such small value of $\hat{p_{j}}$ and tends to be relatively small. The small gradient of negative samples can potentially result in slow or even stagnant optimization, leading to suboptimal performance. We refer to this as gradient vanishing of negative samples under sparse positive feedback. We conclude with the following finding:
Finding 3. When there is sparse positive feedback, the gradients of negative samples vanish since it’s proportional to the estimated positive rate.
(a) BCE vs. BCE with Pairwise
(b) BCE vs. BCE with Listwise
Figure 2. Gradient norm dynamics of negative samples logits in BCE method and classification-ranking combination methods on Criteo dataset in first training epoch.
3.2.2. Gradient of BCE Loss for Positive Sample
@@
$\nabla_{z^{(+)}_{i}}\mathcal{L}_{\text{BCE}}=$ $-\frac{1}{N}\cdot\frac{1}{\sigma(z^{(+)}_{i})}\cdot\sigma(z^{(+)}_{i})(1-\sigma(z^{(+)}_{i}))$
(10) $=$ $-\frac{1}{N}(1-\sigma(z^{(+)}_{i}))=-\frac{1}{N}\cdot({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}1-\hat{p}_{i}}),$
where $N$ denotes the number of all samples. According to Eq. 3.2.2, positive samples satisfy the $\nabla_{z^{(+)}_{i}}\mathcal{L}_{\text{BCE}}\propto 1-\hat{p_{i}}$, which is a rather large value (close to $1$ when $\hat{p_{i}}$ is small) and therefore don’t have gradient vanishing problems as negative samples do. We then examine the gradients of both positive and negative samples on the Combined-Pair method.
3.2.3. Gradient of Combined-Pair for Negative Sample
@@
(11) $\nabla_{z_{j}^{(-)}}\mathcal{L}_{rank}^{CP}=\frac{1}{N_{+}N_{-}}\sum\limits_{i\in N_{+}}\sigma(z_{j}^{(-)}-z_{i}^{(+)}).$
In both our online and offline advertising, when positive feedback is extremely sparse (e.g., the typical CTR for a large-scale online advertising platform is usually less than $2\%$), it is observed that even the estimated values of positive samples tend to be much lower than 0.5. Consequently, the logit of positive samples will also be less than 0. This can result in greater gradients of negative samples in the RankNet Loss, compared to the BCE Loss, as follows:
(12) $\nabla_{z_{j}^{(-)}}\mathcal{L}_{rank}^{CP}=$ $\frac{1}{N_{+}N_{-}}\sum\limits_{i\in N_{+}}\sigma(z_{j}^{(-)}-z_{i}^{(+)})$
$>$ $\frac{1}{N_{-}}\cdot[\frac{1}{N_{+}}N_{+}\cdot\sigma(z_{j}^{(-)})]=\frac{1}{N_{-}}\sigma(z_{j}^{(-)})$
$>$ $\frac{1}{N}\sigma(z_{j}^{(-)})=\nabla_{z_{j}^{(-)}}\mathcal{L}_{\text{BCE}}.$
This indicates that in the sparse positive scenario and for the same negative sample logit, RankNet Loss always has a larger gradient than BCE Loss. Thus, the overall comparison between the BCE method and Combined-Pair method holds:
(13) $\nabla_{z_{j}^{(-)}}\mathcal{L}^{CP}=$ $\alpha\nabla_{z_{j}^{(-)}}\mathcal{L}_{\text{BCE}}+(1-\alpha)\nabla_{z_{j}^{(-)}}\mathcal{L}_{rank}^{CP}$
$>$ $\alpha\nabla_{z_{j}^{(-)}}\mathcal{L}_{\text{BCE}}+(1-\alpha)\nabla_{z_{j}^{(-)}}\mathcal{L}_{\text{BCE}}=\nabla_{z_{j}^{(-)}}\mathcal{L}_{\text{BCE}}.$
@@
Finding 4. When there is sparse positive feedback, Combined-Pair has larger gradients for negative samples than BCE method.
@@ 3.3. Empirical Analysis of Gradient Vanishing
To empirically validate the accuracy of the analysis, we examine the gradient norm dynamics of negative sample logits in the first epoch for the BCE method and Combined-Pair on the Criteo dataset. To simulate the extremely low proportion of positive samples in real-world scenarios, we adjusted the weights to reduce the dataset’s equivalent CTR to approximately 3.3% (detailed setting will be discussed in Section 4.1). As shown in Fig. 2(a), the BCE method (blue line) shows negligible gradient norms of negative samples, while Combined-Pair (red line) shows significantly larger ones. This observation directly supports our analysis in this section.
(a) Bottom Layer of DNN
(b) Bottom Layer of CrossNet
Figure 3. Gradient norm dynamics of DNN and CrossNet in DCN V2 through the training process. pct and avg. are shorts for percentile and average.
Besides, we also care about the optimization issue of trainable parameters, which mainly occurs in trainable parameters rather than outputs (i.e., logits). We hypothesize that the parameters also suffer from gradient vanishing because they obey the chain rule and are affected by gradients of logits.
Thus, we then compare the Combined-Pair and the BCE method on gradient norms of trainable parameters in the bottom layers of DNN and CrossNet. We also report the dynamics of the 90 percentile and average value during training, as shown in Fig. 3. Combined-Pair shows both greater 90 percentiles and average values than the BCE method, persisting throughout the entire training process. This further reflects that Combined-Pair can also effectively alleviate the gradient vanishing problem in learnable parameters.
@@ 4. Experiments
In this section, we conduct empirical experiments to answer the following research questions (RQs): First, we already show that the Combined-Pair outperforms the BCE method since it mitigates the gradient vanishing issue when the positive feedback is sparse. So, if we create artificial datasets with various positive sparsity rates, how would these two methods perform (RQ1)? What’s the trade-off between classification loss and ranking loss (RQ2)? Second, how do other methods with classification-ranking combination loss perform (RQ3)? Last, are there other methods beyond the classification-ranking loss that can also alleviate the gradient vanishing and hence improve classification performance (RQ4)?
@@
(b) LogLoss
Figure 4. Performance evaluation of Combined-Pair and the BCE method under varying positive sparsity rates. Here $\beta_{pos}$ denotes the weights of positive samples.
4.1. Dataset and Experimental Setting
We conducted experiments on the public Criteo dataset (Labs, 2014) [^1] https://github.com/openbenchmark/BARS/tree/main/datasets/Criteo#criteo_x1, a widely used advertising recommendation dataset. It consists of 13 numerical features and 26 categorical features. Specifically, we utilized the criteo_x1 version, where the training, validation, and testing data are divided in a 7:2:1 ratio. Please note that the Criteo dataset’s CTR is 25.6%, which is quite high due to downsampling on negative samples. We introduced a weight $\beta_{pos}$ for the positive samples during model training and evaluation to simulate positive sparsity. We employed DCN V2 (Wang et al., 2021) as the backbone model and utilized the FuxiCTR implementation [^2] https://github.com/xue-pai/FuxiCTR/tree/main/model_zoo/DCNv2 with the same settings as BARS [^3] https://github.com/openbenchmark/BARS/tree/main/ctr_prediction/benchmarks/ DCNv2/DCNv2_criteo_x1. We evaluated the performance regarding two metrics: Logloss (i.e., binary-cross entropy loss) to measure the classification ability and AUC of ROC to measure the ranking ability.
We created several artificial datasets with various sparsity degrees of positive feedback by construction and evaluated both the BCE method and the Combined-Pair. If our hypothesis is true, the datasets with sparser positives suffer more from gradient vanishing. Therefore, the BCE method should perform poorly on these datasets, while the combination loss model should perform much better. Specifically, we build our artificial dataset based on the Criteo dataset and control the sparsity degree of positives by assigning a weight $0<\beta_{pos}\leq 1$ for all its positive samples. This weight is used to down-weight positive samples in the training loss:
(14) $\mathcal{L}_{BCE}=-\frac{1}{N}\sum_{i\in N}[\beta_{pos}\cdot y_{i}\log(\sigma(z_{i}))+(1-y_{i})\log(1-\sigma(z_{i}))].$
For example, by setting $\beta_{pos}=0.1$ for all positive samples, we generate a new dataset with sparsity degree of positive as $25.6\%\cdot 0.1/(25.6\%\cdot 0.1+1-25.6\%)=3.3\%$. We created artificial datasets, with $\beta_{pos}$ equals to $0.8$, $0.6$, $0.4$, $0.2$, $0.1$, respectively. A smaller $\beta_{pos}$ indicates sparser positive feedback. As shown in Fig. 4, we observe that with sparser positive feedback, there is a larger AUC lift (from 0.020% to 0.095%) and LogLoss drop (from 0.045% to 0.168%). This validates our hypothesis that Combined-Pair achieves a larger performance lift than the BCE method when facing sparser positives.
(a) $\beta_{pos}=1.0$
(b) $\beta_{pos}=0.4$
(c) $\beta_{pos}=0.1$
Figure 5. AUC and negative LogLoss of Combined-Pair and BCE method in Criteo test set. $\alpha$ and $1-\alpha$ are the weights of BCE loss and RankNet loss within Combined-Pair, respectively. The scatter with a distinct edge in red represents the pure BCE method.
(a) $\beta_{pos}=1.0$
(b) $\beta_{pos}=0.4$
(c) $\beta_{pos}=0.1$
Figure 6. Training error (i.e., LogLoss) of Combined-Pair and the BCE method on Criteo.
@@ 4.3. RQ2: Trade-off between Classification and Ranking Loss
We aim to investigate the trade-off between the classification and the ranking loss within the combination loss. Specifically, we tune the loss weight $\alpha$ from $1.0$ to $0.1$, and evaluate both negative LogLoss and AUC, as shown in Fig. 5. Please note that the combination loss with $\alpha=1$ equals the BCE method. We have the following observation.
First, when we decrease $\alpha$ from $1.0$, which equals the BCE method, both negative LogLoss and AUC are improved simultaneously, represented by the blue arrow. For example, with $\beta_{pos}=0.1$, decreasing $\alpha$ from $1.0$ to $0.7$ leads to an improvement of negative LogLoss from $0.12152$ to $0.12131$, and AUC from $0.81321$ to $0.81398$. This indicates that the classification and ranking ability can be improved by incorporating ranking loss, with an appropriate weight $\alpha$.
Second, when the $\alpha$ is further increased, both metrics start to deteriorate. However, when the positive is very sparse, e.g., $\beta_{pos}=0.1$, even with a very large weight for the ranking loss, e.g., $0.91-\alpha=$ 0.9, the model’s performance is still better than the BCE method.
@@ 4.4. RQ3: Evaluation of Different Ranking Losses
In this section, we expand our analysis to include various classification-ranking combination methods. We aim to examine whether these methods possess properties similar to the Combined-Pair, thereby enhancing the applicability of our theory to a wider range of scenarios. While Combined-Pair integrates BCE loss with pairwise ranking loss, other methods merge BCE loss with listwise ranking loss. Combined-List (Yan et al., 2022) employs BCE loss in conjunction with the original ListNet loss (Cao et al., 2007). RCR (Bai et al., 2023) combines BCE loss with ListCE loss, a variant of ListNet loss designed to align its minima with that of BCE loss. JRC (Sheng et al., 2023) decoupled the logit into click and non-click logits and proposed a corresponding combination method with listwise-like loss.
These combination methods also incorporate ranking loss with BCE loss. So, can they also alleviate gradient vanishing regarding negative samples? We observe the dynamics of gradient norms of these methods on the Criteo dataset in the first training epoch. Through the analysis shown in Fig. 2(b), we found that their gradient norms are relatively improved compared to the BCE method to varying degrees across different forms, indicating that gradient vanishing alleviation also occurs in combination methods other than Combined-Pair. Among them, since the JRC decouples the logit into click and non-click logits, their gradient magnitudes differ. Thus, JRC has different gradient norms between click/non-click logits. Empirically, the gradient of non-click logit is much larger, and therefore, the role in alleviating gradient vanishing mainly comes from non-click logit.
Table 1. The effectiveness of combining different ranking losses under sparse positive feedback situations (${}^{\triangle}$ represents performance improvement compared to BCE).
@@
AUC 0.81321 ${}^{\triangle}$ ${}^{\triangle}$ ${}^{\triangle}$ ${}^{\triangle}$
LogLoss 0.12152 ${}^{\triangle}$ ${}^{\triangle}$ 0.12152 ${}^{\triangle}$
Furthermore, does their alleviating gradient vanishing also lead to performance improvement? We compare the performance of JRC, Combined-List, and RCR methods on the Criteo dataset with $\beta_{pos}=0.1$. We found they all show improved ranking and classification performance on the Criteo dataset, as shown in Tab. 1. Overall, it can be concluded that other combination methods can also achieve performance improvement by introducing ranking loss to alleviate gradient vanishing in negative samples.

### v2

@@ 3.2. Gradients Analysis
To gain further insights into the reasons behind the aforementioned observations, we conduct a detailed analysis of the gradients in both the BCE and Combined-Pair methods. We begin by examining the gradients of the BCE method. According to the chain rule, the gradients of the parameters in each layer are proportional to the gradients of the logits. Hence, our initial focus is on studying the logit gradients.
3.2.1. Gradients of BCE Loss for Negative Samples
@@
$\nabla_{z^{(-)}_{j}}\mathcal{L}_{\text{BCE}}=$ $\frac{1}{1-\sigma(z^{(-)}_{j})}\cdot\sigma(z^{(-)}_{j})(1-\sigma(z^{(-)}_{j}))$
(10) $=$ $\sigma(z^{(-)}_{j})={\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\hat{p}_{j}}.$
This equation demonstrates that the gradients of negative samples are proportional to its pCTR value, $\hat{p}_{j}$. The expected value of $\hat{p}_{j}$ produced by an unbiased CTR estimation model with BCE loss is close to the underlying global CTR, which equals approximately to the proportion of click samples (i.e., positive feedback) to the total samples. This is because the BCE loss function is scale calibrated (Yan et al., 2022) and its global minima are achieved at $\sigma(z)\rightarrow\mathrm{E}[y|x]$.
When the positive feedback is sparse (e.g., the CTR in our real-world advertising platform is usually less than $2\%$), $\hat{p_{j}}$ becomes a small value. According to Eq. 3.2.1, the gradients of negative samples are proportional to such small value of $\hat{p_{j}}$ and tend to be relatively small. We refer to this as gradient vanishing of negative samples under sparse positive feedback. We conclude this finding as:
Finding 3. When positive feedback is sparse, the gradients of negative samples vanish since they are proportional to the estimated positive rates, which are small in an unbiased estimator.
Figure 2. Gradient norm dynamics of negative samples logits in BCE method and BCE-pairwise ranking combination methods (left) and BCE-listwise ranking combination methods (right) on the Criteo dataset in the first training epoch. All methods set $\alpha=0.5$ in both plots.
3.2.2. Gradients of BCE Loss for Positive Samples
@@
$\nabla_{z^{(+)}_{i}}\mathcal{L}_{\text{BCE}}=$ $-\frac{1}{\sigma(z^{(+)}_{i})}\cdot\sigma(z^{(+)}_{i})(1-\sigma(z^{(+)}_{i}))$
(11) $=$ $-(1-\sigma(z^{(+)}_{i}))=-({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1-\hat{p}_{i}}).$
According to Eq. 3.2.2, positive samples satisfy the $\nabla_{z^{(+)}_{i}}\mathcal{L}_{\text{BCE}}\propto 1-\hat{p_{i}}$, which is a relatively large value (close to $1$ when $\hat{p_{i}}$ is small) and therefore don’t have gradient vanishing problems as negative samples do.
3.2.3. Gradients of Combined-Pair for Negative Samples
@@
(12) $\nabla_{z_{j}^{(-)}}\mathcal{L}_{\text{Rank}}^{\text{CP}}=\frac{1}{N_{+}}\sum_{i=1}^{N_{+}}\sigma(z_{j}^{(-)}-z_{i}^{(+)}).$
In both our online and offline advertising, when positive feedback is extremely sparse, it is observed that even the estimated values of positive samples tend to be much lower than 0.5. Consequently, the logit of positive samples $z_{i}^{(+)}$ is less than 0. This can result in greater gradients of negative samples in the RankNet Loss, compared to the BCE Loss, as follows:
(13) $\nabla_{z_{j}^{(-)}}\mathcal{L}_{\text{Rank}}^{\text{CP}}=$ $\frac{1}{N_{+}}\sum_{i=1}^{N_{+}}\sigma(z_{j}^{(-)}-z_{i}^{(+)})$
$>$ $\frac{1}{N_{+}}\cdot N_{+}\cdot\sigma(z_{j}^{(-)})$
$=$ $\sigma(z_{j}^{(-)})=\nabla_{z_{j}^{(-)}}\mathcal{L}_{\text{BCE}}.$
This indicates that in the sparse positive scenario and for the same negative sample logit, RankNet Loss may have larger gradients than BCE Loss. Thus, the following inequation between the BCE method and the Combined-Pair method holds:
(14) $\nabla_{z_{j}^{(-)}}\mathcal{L}^{\text{CP}}=$ $\alpha\nabla_{z_{j}^{(-)}}\mathcal{L}_{\text{BCE}}+(1-\alpha)\nabla_{z_{j}^{(-)}}\mathcal{L}_{\text{Rank}}^{\text{CP}}$
$>$ $\alpha\nabla_{z_{j}^{(-)}}\mathcal{L}_{\text{BCE}}+(1-\alpha)\nabla_{z_{j}^{(-)}}\mathcal{L}_{\text{BCE}}$
$=$ $\nabla_{z_{j}^{(-)}}\mathcal{L}_{\text{BCE}}.$
@@
Finding 4. When positive feedback is sparse, Combined-Pair has larger gradients for negative samples than the BCE method.
@@ 3.3. Empirical Analysis of Gradient Vanishing
To empirically validate the analysis, we examine the dynamics of gradient norms for negative sample logits in the first epoch of training. To simulate the low proportion of positive samples often encountered in real-world scenarios, we adjust the positive sample rates, denoted as $\beta_{\text{pos}}$, to achieve an equivalent Click-Through Rate (CTR) of 3.3% for the dataset. Further details regarding this adjustment will be discussed in Sec. 4.1.
As depicted in Fig. 2, we observe that the BCE method (the blue line) exhibits negligible gradient norms for negative samples. In contrast, the Combined-Pair method (the red line) demonstrates significantly larger gradient norms for negative samples. This empirical observation aligns with our previous analysis of the gradients.
Figure 3. Gradient norm dynamics of DNN (left) and CrossNet (right) in DCN V2 through the training process. pct and avg. are shorts for percentile and average.
We further investigate the optimization procedure of the trainable parameters in the entire model architecture. Specifically, we compare the Combined-Pair method with the BCE method by examining the gradient norms of the trainable parameters in the bottom layers of the Deep Neural Network (DNN) and CrossNet in DCN V2. To provide a comprehensive understanding of the optimization dynamics, we report the dynamics of the 90th percentile and average values of the gradient norms during the training process. These results are depicted in Fig. 3. Remarkably, we observe that the Combined-Pair method consistently achieves higher values on both metrics compared to the BCE method. This difference persists throughout the entire training process. This finding further validates that the Combined-Pair method effectively alleviates the issue of gradient vanishing in the learnable parameters.
@@ 4. Experiments
In this section, we conduct empirical experiments to answer the following research questions (RQs): If the sparse positive rate is the main cause of the gradient vanishing of negative samples, how would models perform under various sparsity of positives? What’s the trade-off between classification loss and ranking loss (RQ2)? How do the other methods with classification-ranking combination losses perform (RQ3)? Last, can our perspective be extended to methods beyond the classification-ranking combination loss (RQ4)?
4.1. Dataset and Experimental Setting
We conducted experiments on the public Criteo dataset [^1] https://github.com/reczoo/Datasets/tree/main/Criteo/Criteo_x1 (Labs, 2014), a widely used advertising recommendation dataset. It consists of 13 numerical features and 26 categorical features. Specifically, we utilized the criteo_x1 version, where the training, validation, and testing data are divided in a 7:2:1 ratio. We employed DCN V2 (Wang et al., 2021) as the backbone model and utilized the FuxiCTR (Zhu et al., 2021) implementation [^2] https://github.com/reczoo/FuxiCTR/tree/main/model_zoo/DCNv2 with the same settings as BARS (Zhu et al., 2022) [^3] https://github.com/reczoo/BARS/tree/main/ranking/ctr/DCNv2/DCNv2_criteo_x1. We evaluated the performance regarding two metrics: BCE loss (i.e., binary-cross entropy loss) to measure the classification ability and AUC to measure the ranking ability.
Specifically, we create artificial datasets based on the Criteo dataset and control the sparsity degree of positives by assigning a weight $0<\beta_{\text{pos}}\leq 1$ for all its positive samples. This weight is used to down-weight positive samples in the training loss:
(15) $\mathcal{L}_{BCE}=-\frac{1}{N}\sum_{i=1}^{N}[\beta_{\text{pos}}\cdot y_{i}\log(\sigma(z_{i}))+(1-y_{i})\log(1-\sigma(z_{i}))].$
For example, by setting $\beta_{\text{pos}}=0.1$ for all positive samples, we generate a new dataset with sparsity degree of positive as $\frac{25.6\%\times 0.1}{25.6\%\times 0.1+1-25.6\%}=3.3\%$.
To validate our hypothesis, we generated multiple artificial datasets with varying degrees of sparsity in positive feedback by adjusting the values of $\beta_{\text{pos}}$. We then evaluated the performance of both the BCE method and the Combined-Pair method on these datasets. If our hypothesis holds true, datasets with sparser positive feedback should be more susceptible to the issue of gradient vanishing. Consequently, we would expect the BCE method to exhibit poorer performance on these datasets, while the Combined-Pair method should demonstrate significantly better performance, resulting in a larger performance gap compared to the BCE method.
@@ 4.2. RQ1: Performance Evaluation with various Positive Sparsity Rates
(b) BCE loss
Figure 4. Performance evaluation of Combined-Pair and the BCE method under varying positive sparsity rates. Here $\beta_{\text{pos}}$ denotes the weights of positive samples.
In particular, we created artificial datasets with $\beta_{\text{pos}}$ equals to $0.8$, $0.6$, $0.4$, $0.2$, $0.1$, respectively. A smaller $\beta_{\text{pos}}$ indicates sparser positive feedback. As shown in Fig. 4, we observe that the Combined-Pair always gets better AUC and BCE loss at all $\beta_{\text{pos}}$. Especially, from $\beta_{\text{pos}}=0.6$ to $\beta_{\text{pos}}=0.1$, the sparser positive rates (i.e., smaller $\beta_{\text{pos}}$), the larger AUC lift (from 0.020% to 0.095%) and BCE loss drop (from 0.045% to 0.168%) between the Combined-Pair and the BCE method. This validates our hypothesis that Combined-Pair achieves a larger performance lift than the BCE method when sparser positives reach a sparsity threshold (here $\beta_{\text{pos}}=0.6$).
@@ 4.3. RQ2: Trade-off between Classification and Ranking Loss
Our aim to examine the trade-off between the classification and ranking loss components within the Combined-Pair loss. To achieve this, we vary the loss weight parameter, denoted as $\alpha$, in the Combined-Pair loss from $1.0$ to $0.1$. We evaluate the negative BCE loss and the Area Under the Curve (AUC) as performance metrics, as depicted in Fig. 5. Based on our observations, we note the following:
First, starting with $\alpha=1.0$, which corresponds to the Binary Cross-Entropy (BCE) method (represented by the red diamond), we observe that decreasing $\alpha$, i.e., reducing the weight of the BCE loss while increasing the weight for the ranking loss, leads to simultaneous improvements in both the negative BCE loss and AUC. This trend is represented by the orange arrow in the figure. For instance, with $\beta_{\text{pos}}=0.1$, reducing $\alpha$ from 1.0 to 0.7 results in a decrease in negative BCE loss from 0.1215 to 0.1213, and an increase in AUC from 0.8132 to 0.8139. This suggests that the classification and ranking abilities can be improved monotonically by decreasing $\alpha$ up to a certain threshold.
Second, we observe that both metrics deteriorate when $\alpha$ is further increased. In other words, as the ranking loss becomes more dominant in the combination loss beyond a certain threshold, both the classification and ranking abilities deteriorate monotonically (as shown by the blue arrow).
However, when the positive feedback is very sparse (e.g., $\beta_{\text{pos}}=0.1$), even with a very large weight for the ranking loss (e.g., $0.91-\alpha=$ 0.9), the model’s performance remains superior to the BCE method.
Figure 5. AUC and negative BCE loss of Combined-Pair and BCE method in Criteo test set. $\alpha$ and $1-\alpha$ are the weights of BCE loss and RankNet loss within Combined-Pair, respectively. The diamond in red represents the BCE method, i.e., $\alpha=1$. We shows results with $\alpha$ ranging between $[0.1,1]$
@@ 4.4. RQ3: Evaluation of Different Ranking Losses
In this section, we expand our analysis to include other classification-ranking combination methods beyond Combined-Pair. We aim to examine whether these methods possess properties similar to the Combined-Pair, thereby enhancing the applicability of our theory to a broader range of methods.
While Combined-Pair integrates BCE loss with pairwise ranking loss, other methods combine BCE loss with listwise ranking loss. For example, Combined-List (Yan et al., 2022) employs BCE loss in conjunction with the original ListNet loss (Cao et al., 2007). RCR (Bai et al., 2023) combines BCE loss with ListCE loss, a variant of ListNet loss designed to align its minima with that of BCE loss. JRC (Sheng et al., 2023) decoupled the logit into click and non-click logits and proposed a corresponding combination method with listwise-like loss.
These combination methods also incorporate ranking loss with BCE loss. So, can they also alleviate gradient vanishing regarding negative samples? We monitor the gradient norm dynamics of negative samples for these methods on the Criteo dataset in the first training epoch. Through the analysis shown in Fig. 2, we found that their gradient norms are relatively improved compared to the BCE method to varying degrees across different forms, indicating that gradient vanishing is also alleviated in these methods. Among them, JRC decouples the logit into click and non-click logit, and the gradient vanishing is mainly mitigated on the non-click logit.
Table 1. The performance of combining different ranking losses under sparse positive feedback situations. The $\uparrow$ and $\downarrow$ represent increasing in AUC and decreasing in BCE loss compared to the BCE method, respectively.
@@
[…]
BCE $\downarrow$ 0.12152 0.12131 0.12146 0.12152 0.12141
We then compare the performance of JRC, Combined-List, and RCR methods on the Criteo dataset with $\beta_{\text{pos}}=0.1$. We found they all show improved ranking and classification performance, as shown in Tab. 1. Overall, it can be concluded that these combination methods can also achieve performance improvement by introducing ranking loss to alleviate gradient vanishing of negative samples.
(a) $\beta_{\text{pos}}=1.0$
(b) $\beta_{\text{pos}}=0.4$
(c) $\beta_{\text{pos}}=0.1$
Figure 6. Training error (i.e., BCE loss) of the BCE method and Combined-Pair with various $\beta_{\text{pos}}$ on Criteo.